# MPG-SAM 2: Adapting SAM 2 with Mask Priors and Global Context for Referring Video Object Segmentation

### v2

@@ Abstract
Referring video object segmentation (RVOS) aims to segment objects in a video according to textual descriptions, which requires the integration of multimodal information and temporal dynamics perception. The Segment Anything Model 2 (SAM 2) has shown great effectiveness across various video segmentation tasks. However, its application to offline RVOS is challenged by the translation of the text into effective prompts and a lack of global context awareness. In this paper, we propose a novel RVOS framework, termed MPG-SAM 2, to address these challenges. Specifically, MPG-SAM 2 employs a multimodal encoder to jointly encode video and textual features, generating semantically aligned video and text embeddings along with multimodal class tokens. A mask prior generator is devised to utilize the video embeddings and class tokens to create pseudo masks of target objects and global context. These masks are fed into the prompt encoder as dense prompts, along with multimodal class tokens as sparse prompts to generate accurate prompts for SAM 2. To provide the online SAM 2 with a global view, we propose a hierarchical global-historical aggregator, which allows SAM 2 to aggregate global and historical information of target objects at both pixel and object levels, enhancing the target representation and temporal consistency. Extensive experiments on several RVOS benchmarks demonstrate the superiority of MPG-SAM 2 and the effectiveness of the proposed modules.
@@ 2 Related work
Referring Video Object Segmentation. RVOS attracts extensive interest as it bridges visual and linguistic domains. Early methods, such as RefVOS [2], consider RVOS as an expansion of referring image segmentation (RIS) to the video domain. URVOS [35] advances this by combining RIS and semi-supervised video object segmentation within a unified framework using attention mechanisms. In subsequent studies [28, 38, 21, 12, 40], researchers emphasize cross-modal interaction, further enhancing RVOS performance. However, despite notable gains, the computational expense and complexity of multi-stage pipelines limit practical feasibility. In response, query-based Transformer architectures offer efficient solutions with simplified yet robust frameworks. Notably, MTTR [3] and ReferFormer [42] pioneer the use of the DETR series [5, 53] in RVOS, introducing novel multimodal interaction mechanisms. Recent methods [26, 29, 36, 20] refine these Transformer architectures through advanced temporal and multimodal feature integration techniques. For instance, SgMg [29] enhances ReferFormer by replacing dynamic convolution with a segmentation optimizer and uses spectral information to guide visual feature fusion. SOC [26] implements layered temporal modeling across video and object levels, achieving early multimodal fusion. More recently, Losh [49] introduces a joint prediction network for short and long sentences, reducing the over-influence of action and relational cues in segmentation. VD-IT [54] explores the potential of text-to-video diffusion models in the RVOS task, leveraging the inherent rich semantics and coherent temporal correspondences of video generation models to ensure temporal instance consistency. To tackle longer sequences and complex scenarios in the MeViS [11] dataset, DsHmp [14] proposes a static text decoupling strategy, enhancing the temporal understanding of static and dynamic content at both frame and object levels, thereby capturing both short-term and long-term motion insights.
@@ 3.4 Hierarchical Global-Historical Aggregator
Unlike VOS that primarily relies on past information, RVOS emphasizes the effective utilization of both temporal and textual cues. Inspired by [7], we propose a hierarchical global-historical aggregator that synergistically combines SAM 2’s memory mechanism with multi-level temporal modeling, enabling comprehensive integration of global context and historical segmentation results at both pixel-wise and object-centric levels. As shown in Fig. 3, the hierarchical global-historical aggregator comprises two components: the pixel-level fusion module and the object-level fusion module.

### v3

@@ Abstract
Referring video object segmentation (RVOS) aims to segment objects in a video according to textual descriptions, which requires the integration of multimodal information and temporal dynamics perception. The Segment Anything Model 2 (SAM 2) has shown great effectiveness across various video segmentation tasks. However, its application to offline RVOS is challenged by the translation of the text into effective prompts and a lack of global context awareness. In this paper, we propose a novel RVOS framework, termed MPG-SAM 2, to address these challenges. Specifically, MPG-SAM 2 employs a multimodal encoder to jointly encode video and textual features, generating semantically aligned video and text embeddings along with multimodal class tokens. A mask prior generator is devised to utilize the video embeddings and class tokens to create pseudo masks of target objects and global context. These masks are fed into the prompt encoder as dense prompts, along with multimodal class tokens as sparse prompts to generate accurate prompts for SAM 2. To provide the online SAM 2 with a global view, we propose a hierarchical global-historical aggregator, which allows SAM 2 to aggregate global and historical information of target objects at both pixel and object levels, enhancing the target representation and temporal consistency. Extensive experiments on several RVOS benchmarks demonstrate the superiority of MPG-SAM 2 and the effectiveness of the proposed modules. The code is available at https://github.com/rongfu-dsb/MPG-SAM2.
@@ 2 Related work
Referring Video Object Segmentation. RVOS attracts extensive interest as it bridges visual and linguistic domains. Early methods, such as RefVOS [2], consider RVOS as an expansion of referring image segmentation (RIS) to the video domain. URVOS [37] advances this by combining RIS and semi-supervised video object segmentation within a unified framework using attention mechanisms. In subsequent studies [29, 40, 22, 12, 43], researchers emphasize cross-modal interaction [13, 46], further enhancing RVOS performance. However, despite notable gains, the computational expense and complexity of multi-stage pipelines limit practical feasibility. In response, query-based Transformer architectures offer efficient solutions with simplified yet robust frameworks. Notably, MTTR [3] and ReferFormer [45] pioneer the use of the DETR series [5, 57] in RVOS, introducing novel multimodal interaction mechanisms. Recent methods [27, 30, 38, 21, 33] refine these Transformer architectures through advanced temporal and multimodal feature integration techniques. For instance, SgMg [30] enhances ReferFormer by replacing dynamic convolution with a segmentation optimizer and uses spectral information to guide visual feature fusion. SOC [27] implements layered temporal modeling across video and object levels, achieving early multimodal fusion. More recently, Losh [53] introduces a joint prediction network for short and long sentences, reducing the over-influence of action and relational cues in segmentation. VD-IT [58] explores the potential of text-to-video diffusion models in the RVOS task, leveraging the inherent rich semantics and coherent temporal correspondences of video generation models to ensure temporal instance consistency. To tackle longer sequences and complex scenarios in the MeViS [11] dataset, DsHmp [15] proposes a static text decoupling strategy, enhancing the temporal understanding of static and dynamic content at both frame and object levels, thereby capturing both short-term and long-term motion insights.
@@ 3.4 Hierarchical Global-Historical Aggregator
Unlike VOS that primarily relies on past information, RVOS emphasizes the effective utilization of both temporal and textual cues. Inspired by [7], we propose a hierarchical global-historical aggregator that synergistically combines SAM 2’s memory mechanism with multi-level temporal modeling, enabling comprehensive integration of global context and historical segmentation results at both pixel-wise and object-centric levels. As shown in Tab. 7, the hierarchical global-historical aggregator comprises two components: the pixel-level fusion module and the object-level fusion module.
@@ 5 Conclusion
Supplementary Material
Appendix A Additional Experimental Studies
Mask-text Similarity Loss. In this part, We validate the generalizability of the mask-text similarity loss function by conducting enhancement experiments with this function on several previous RVOS methods, including ReferFormer [45] and SgMg [30]. Meanwhile, we also validate the impact of this function on the performance of MPG-SAM 2. The experimental results, presented in Tab. 5, indicate performance improvements across all methods, confirming the effectiveness of the proposed similarity function in RVOS tasks.
Model Parameters. In this section, we analyze the parameter count of our model. We supplement our study with a set of low-configuration experiments on the Ref-YouTube-VOS [37] dataset, using SAM 2-Hiera-Large [36] and BEiT-3-Base [41] as initialization parameters, called MPG-SAM 2-Tiny. The experimental results and parameter counts are presented in Tab. 6. Compared to previous methods with relatively small parameter sizes, such as ReferFormer [45] and SgMg [30], our low-configuration model MPG-SAM 2-Tiny exhibits a slightly larger parameter count but achieves a substantial performance gain. Furthermore, although methods like VISA [49] and HyperSeg [42] also employ vision-language models and are trained on additional datasets, in contrast, our full-configuration model MPG-SAM 2, which is not trained on any additional datasets, demonstrates superior performance with a smaller model size. This highlights the effectiveness and efficiency of our approach.
Hierarchical Global-Historical Aggregator. In this division, we perform a more detailed component analysis to evaluate the effectiveness of each part within the hierarchical global-historical aggregator (HGA) on the Ref-YouTube-VOS [37] dataset. The experimental results are presented in Tab. 7. The initial setup involves the MPG-SAM 2 only using the mask prior generator, which achieves 71.9% $\mathcal{J}\&\mathcal{F}$. When incorporating the global enhancement and local enhancement of HGA’s object-level fusion part separately, the model reaches $\mathcal{J}\&\mathcal{F}$ scores of 72.2% and 72.4%, respectively, resulting in gains of 0.3% and 0.5% $\mathcal{J}\&\mathcal{F}$. When applying both components of the object-level fusion part to the initial model simultaneously, the model attains 72.8% $\mathcal{J}\&\mathcal{F}$, indicating a 0.9% $\mathcal{J}\&\mathcal{F}$ score improvement compared to the initial setup. Moreover, the inclusion of HGA’s pixel-level global fusion part on the initial setup obtains a $\mathcal{J}\&\mathcal{F}$ score of 73.1%. Finally, when all parts of HGA are employed, the full model gains the highest performance of 73.9% $\mathcal{J}\&\mathcal{F}$.
Table 5: Generalizability of the mask-text similarity loss.
[…]
ReferFormer [45] ResNet-50 55.6 54.8 56.5
ReferFormer [45] + $\mathcal{L}_{sim}$ ResNet-50 56.4 55.4 57.5
SgMg [30] Video-Swin-T 62.0 60.4 63.5
SgMg [30] + $\mathcal{L}_{sim}$ Video-Swin-T 62.8 61.3 64.3
MPG-SAM 2 - $\mathcal{L}_{sim}$ Hiera-L 73.2 71.2 75.2
MPG-SAM 2 Hiera-L 73.9 71.7 76.1
Table 6: Model parameter analysis on Ref-YouTube-VOS dataset. Our model strikes a balance between the number of parameters and performance, demonstrating clear advantages. The best results are high-lighted in bold, and the second best results are underlined.
[…]
ReferFormer [45] CVPR’22 0.24B 62.9 61.3 64.6
SgMg [30] ICCV’23 0.24B 65.7 63.9 67.4
VISA [49] ECCV’24 13B 63.0 61.4 64.7
HyperSeg [42] Arxiv’24 3B 68.5 - -
MPG-SAM 2-Tiny - 0.46B 69.9 68.0 71.8
MPG-SAM 2 - 0.92B 73.9 71.7 76.1
We also analyze the effect of the number of the pixel-level fusion layer $N_{p}$ and object-level fusion layer $N_{o}$ on the model’s performance. For $N_{p}$ and $N_{o}$, we design experiments with 1, 2, and 3 layers, and the results are shown in Tab. 8. The results show that single-layer fusion modules are sufficient to effectively enhance global and historical information at both the pixel level and object level. However, increasing the number of layers introduces redundant information, which may impair the segmentation process for the current frame. As a result, we set both the $N_{p}$ and $N_{o}$ to 1 to ensure optimal model performance.
Table 7: The ablation experiments of HGA components, where OGF represents the object-level global fusion part, OLF denotes the object-level local fusion part and PGF refers to the pixel-level global fusion part of HGA.
[…]
MPG-SAM 2 71.9 69.8 73.9
MPG-SAM 2 72.2 70.1 74.3
MPG-SAM 2 72.4 70.1 74.6
MPG-SAM 2 72.8 70.7 75.0
MPG-SAM 2 73.1 71.0 75.3
MPG-SAM 2 73.9 71.7 76.1
Table 8: Performance analysis of the hierarchical global-historical aggregator with varying numbers of layer.
[…]
MPG-SAM 2 $N_{p}=1$ 73.9 71.7 76.1
MPG-SAM 2 $N_{p}=2$ 73.6 71.6 75.7
MPG-SAM 2 $N_{p}=3$ 73.0 70.9 75.2
MPG-SAM 2 $N_{o}=1$ 73.9 71.7 76.1
MPG-SAM 2 $N_{o}=2$ 73.4 71.2 75.6
MPG-SAM 2 $N_{o}=3$ 72.9 70.7 75.1
Figure 5: Additional visualization results on several datasets. (a), (b) MeViS, (c), (d) Ref-YouTube-VOS.
Appendix B Additional Visualization Results
In this section, we present additional visualization results of our MPG-SAM 2 on Ref-YouTube-VOS [37] dataset and MeViS [11] dataset. The visualizations, shown in Fig. 5, highlight target objects covered by blue masks.
For the MeViS [11] dataset, we evaluate the model’s segmentation performance in challenging scenarios involving object occlusion and multiple referential targets. As shown in Fig. 5 (a), the target object, a tiger, is initially occluded by an enclosure in the first three frames and becomes visible in the subsequent four frames. The model effectively detects the absence of the target in the occluded frames and accurately segments it once it appears. Additionally, Fig. 5 (b) presents a scene with three distinct referential targets, all of which are precisely segmented by the model without any omissions across frames.
For the Ref-YouTube-VOS [37] dataset, we select several complex scenarios featuring multiple similar objects to evaluate the model’s ability to distinguish between such targets. As shown in Fig. 5 (c), the model accurately segments the specific sheep matching the language description from a group of visually similar sheep. In Fig. 5 (d), the model successfully disregards the interference caused by shadows and accurately segments the kangaroo positioned on the left. These findings highlight the robustness and effectiveness of our model in tackling diverse and challenging scenarios.